OpenStudy (anonymous):

Can you find the center and the radius of the circle from x2 + y2 -8x +4y =5

myininaya (myininaya):

\[(x^2-8x)+(y^2+4y)=5\] \[(x^2-8x+(\frac{8}{2})^2)+(y^2+4y+(\frac{4}{2})^2)=5+(\frac{8}{2})^2+(\frac{4}{2})^2\]

OpenStudy (anonymous):

I think C=(4,-2) and radius r=5.

myininaya (myininaya):

\[(x-4)^2+(y+2)^2=5+16+4\]

myininaya (myininaya):

\[(x-4)^2+(y+2)^2=25\]

myininaya (myininaya):

yes thats right emurr

OpenStudy (anonymous):

thanks myininaya :)

myininaya (myininaya):

np